Most of the light form a sodium vapor lamp has a wavelength of 598nm. What is the frequency of this radiation?
1
Expert's answer
2020-11-09T13:53:34-0500

Formula : Frequency = "\frac {C}{\lambda}"

here C = "3\times 10^8 m/s." and "\lambda = 598\times 10^-9 m"

Hence Frequency = "\frac {3\times 10^8}{598\times 10^-9}" = "5.02\times 10^(14) Hz = 5.02\times 10" ^(14) Hz. Answer.